### v2

@@ 5.2.1 Communities comparison
Figure 9: Bar charts showing the number of lost, common, and gained communities switching from INDI to UNFL (sum) and MULTI.
(a) UNFL (sum) vs. INDI.
(b) MULTI vs. INDI.
@@
Figure 10: Bar charts showing the number of lost, common, and gained nodes switching from INDI to UNFL (sum) and MULTI.
@@
We categorize communities as lost, common, or gained when exploiting UNFL and MULTI with respect to INDI. Figure 9 presents bar charts illustrating the number of communities across these categories, and Figure 10 the labelling for the nodes.
@@
Multimodal approaches comparison. Figure 9 shows that more communities are lost when switching to UNFL than to MULTI. While this may seem counterintuitive given that MULTI is more restrictive, this result underscores an important limitation of network flattening. Indeed, the process of flattening can lead to the loss of crucial network information and the introduction of noise in the network topology. This highlights the value of the MULTI approach, which preserves the structure and information inherent in individual layers, unlike UNFL. These findings also call into question the efficacy of the commonly used UNFL strategy, which is prevalent in the literature [31, 26, 52, 2, 30].
(c) UNFL vs. HST
(d) MULTI vs. HST
(e) UNFL vs. MEN
(f) MULTI vs. MEN
(g) UNFL vs. RPL
(h) MULTI vs. RPL
(i) UNFL vs. RTW
(j) MULTI vs. RTW
(k) UNFL vs. URL
(l) MULTI vs. URL
Figure 11: Lost, common, gained nodes characterization for UNFL and MULTI. We use the notation ns and * to indicate statistical significance levels. Specifically, ns denotes that there is no significant difference between the groups, with a p-value in the range $p\geq 0.05$. For statistically significant results, we used the following annotations based on p-value ranges: *: $p\leq 0.05$; **: $p\leq 0.01$; ***: $p\leq 0.001$.
@@
Union flattening. Figure 9(a) provides insights on the results of UNFL with respect to INDI. Hence, UNFL exhibits similar patterns with blocks of common and gained communities and relatively few lost ones with respect to HST, MEN, and RTW. However, it shows a higher number of lost and gained communities compared with URL, suggesting that its information is poorly captured in the flattened network. Finally, UNFL has two communities in common with RPL and no lost communities but a significant number of gained ones, indicating that the detected communities differ substantially from those in the monomodal approach.
@@
Multiplex community detection. Figure 9(b) demonstrates that MULTI has very few lost communities with respect to all monomodal approaches, with the exception of the URL. The transition from the co-reply layer to the multimodal approach reveals a higher number of gained communities, while other co-actions primarily show common communities and fewer gained ones. This suggests that communities identified using the multimodal MULTI approach retain much of the information from the monomodal approach while incorporating additional insights from the interplay between layers.
@@
Discussion. This analysis highlights the effectiveness of MULTI in preserving existing information while uncovering new coordinated communities. Meanwhile, the UNFL is more prone to losing communities, further underscoring its limitations. The results suggest that MULTI is more effective in capturing the underlying structure and information of the network, providing a more comprehensive view of coordinated behavior. The node-level analysis reveals similar trends, with MULTI outperforming the UNFL in retaining and identifying nodes. This comparison underscores the importance of selecting an appropriate multimodal approach to effectively capture the complexity and richness of coordinated behavior in multiplex networks.
@@ 5.2.2 Multimodal communities characterization
To assess the effectiveness of multimodal approaches in identifying central nodes, we analyze the lost, common, and gained nodes defined in Section 3.3.3 and shown in Figure 10. This evaluation aims to determine whether multimodal methods retain or lose key influential nodes. Losing peripheral nodes is not critical, but missing highly connected or bridging nodes would indicate a limitation. Conversely, successfully preserving or discovering new influential nodes would highlight the advantage of multimodality in capturing a more complete network structure. To conduct this analysis, we compute key centrality and structural metrics for each node including degree centrality, eigenvector centrality, local clustering coefficient, and PageRank [75, 76]. The comparison is performed for both UNFL and MULTI. Figure 11 illustrates the distribution of these metrics across lost, common, and gained nodes for each co-action. We applied the Brunner-Munzel test [77] to compare distributions across the lost, common, and gained node subsets for each metric. Unlike the Mann-Whitney U test [78], the Brunner-Munzel test does not assume equal variances or distribution shapes, making it more robust to ties, outliers, and variance differences, and thus more suitable for our analysis. The comparison highlights the performance of both the UNFL and MULTI, offering insights into their ability to retain or identify influential nodes.
@@
Union flattening. Lost and gained nodes are characterized by a higher degree centrality and eigenvector centrality switching from HST, MEN, RTW, and URL to UNFL, with respect to the common nodes. The unique exception is RPL, where the common nodes have lower values compared to gained and lost nodes. Instead, local clustering coefficient distribution is more homogeneous among the different co-actions. Finally, the page rank is higher for the lost and gained nodes, with respect to the common nodes, for HST, MEN, and RTW. Instead, for RPL and URL, the common nodes have higher values, but with similar distributions to the lost and gained ones.
@@
Therefore, in case of HST, MEN, RTW, and URL, UNFL seems to lose the most important nodes, which are the most central and influential in the network, but at the same time it is able to detect new nodes that are not detected by the monomodal approach and that are as well central and influential. Instead, RPL shows a different pattern, where the common nodes are the most central and influential, while the lost and gained nodes are less central and influential, suggesting that both the monomodal approach and the UNFL are able to capture the most important nodes in the network.
@@
Multiplex community detection. Observing MULTI, the common nodes have higher values of degree centrality and eigenvector centrality, with respect to the lost and gained nodes, for all the co-actions, with the exception of URL, where the lost nodes have higher values. Local clustering coefficient shows similar average for the lost, common, and gained nodes, for MEN, but with a higher variance for the lost and gained nodes. HST, RPL, and RTW show higher values for the common nodes, with respect to the lost and gained nodes. Finally, page rank is on average similar for all the co-actions and for the lost, common, and gained nodes, but with a higher variance for the common nodes.
@@
Discussion. From the perspective of node centrality, the MULTI approach yields results that are generally consistent with those of the individual layers. Previous analyses showed that it retains most of the key nodes and communities from each modality, and although it tends to detect additional ones, these are typically less central. In contrast, UNFL exhibits the opposite behavior: it both gains and loses a notable number of nodes and communities, but crucially, it often discards highly central ones while still identifying new, influential actors. This suggests a significant misalignment between the results of UNFL and those of the individual layers. As such, these findings may raise concerns about the reliability of UNFL, especially in terms of preserving meaningful structural information.
@@ 6 Discussion and conclusions
We explored the role of multimodality in the detection and analysis of coordinated behavior, addressing its challenges and opportunities. To tackle this complexity, we introduced a novel methodology that integrates multiple types of user actions into a unified analytical framework for detecting multimodal coordinated behavior. By comparing monomodal approaches with flattened representations, and the multiplex community detection, our analysis highlighted the strengths and weaknesses of each approach, offering valuable insights into the detection and characterization of coordination in social media.
@@ 6.1 RQ1: Monomodal contributions
The second research question examines the contribution of multimodal approaches compared to monomodal ones in analyzing coordinated behavior. Appendix A shows that UNFL implemented with different flattening strategies produces similar outcomes, suggesting that the specific method of flattening has limited influence on the results. However, Section 5.2.1 demonstrates that the act of network flattening itself significantly impacts the detection of coordinated communities. In particular, the UNFL approach shares only a subset of communities with RTW, MEN, and HST, while failing to capture nearly all of those identified by URL. This indicates a substantial loss of information contributed by each monomodal layer. In contrast, the MULTI approach shows greater robustness: it retains almost all communities detected by monomodal methods and also uncovers additional coordinated structures. These findings raise concerns about the use of network flattening in the literature and advocate for the adoption of more expressive, though complex, multimodal methods. Section 5.2.2 further supports the advantage of MULTI, highlighting its ability to preserve influential nodes while also identifying new ones. By comparison, UNFL tends to lose highly central nodes, although it occasionally detects new, equally important ones. Overall, MULTI proves to be a more reliable approach for capturing both the breadth and depth of coordinated behavior.
@@ 6.3 Limitations and future works
This work presents the first comparative analysis of methods to opeartionalize the detection off multimodal coordinated behavior. Indeed, we define multiple operationalizations of multimodality, assessing the contributions of individual monomodal approaches, and evaluating the benefits of a multimodal framework over monomodal ones. The primary limitation of this study is that the analysis was conducted on a single dataset. As a result, the generalizability of our findings remains uncertain, and the conclusions drawn may be specific to the characteristics and dynamics of the observed case [79]. Future work will need to validate the proposed methodology across multiple datasets, ideally covering diverse social phenomena and contexts, to assess the robustness and consistency of our conclusions. Such efforts will be crucial to determine whether the observed patterns hold more broadly and to establish the general applicability of our multimodal framework. Given the lack of prior literature on multimodal coordination, we addressed several foundational research questions, necessarily leaving out additional operationalizations of multimodality. For instance, exploring alternative community detection algorithms, such as multilayer clique percolation [80] and ABACUS [81], could introduce new levels of multimodality within the intrinsic trade-off of coordination analysis (see Figure 2). Furthermore, future work could refine the integration across layers by systematically investigating the $\omega$ parameter of the Generalized Louvain algorithm, which controls the strength of inter-layer connections between the same user across modalities. Our current study focuses on five common co-actions, but the framework could be expanded to incorporate additional modalities, such as multimedia content sharing, which plays a central role on many social media platforms.

### v3

@@ 5.2.1 Communities comparison
Figure 10: Bar charts showing in the number of lost, common, and gained communities switching from INDE to UNFL (sum) and MULTI (in IORussia and UK dataset).
@@
Figure 11: Bar charts showing the number of lost, common, and gained nodes switching from INDE to UNFL (sum) and MULTI (in IORussia and UK dataset).
@@
We categorize communities as lost, common, or gained when exploiting UNFL and MULTI with respect to INDE. We report the results for UK and IORussia with Louvain approach in Figure 10, which presents bar charts illustrating the number of communities across these categories, and in Figure 11 that shows the labelling for the nodes.
@@
Multimodal approaches comparison. Both for UK and IORussia, Figure 10 shows that more communities are lost when switching to UNFL than to MULTI. While this may seem counterintuitive given that MULTI is more restrictive, this result underscores an important limitation of network flattening. Indeed, the process of flattening can lead to the loss of crucial network information and the introduction of noise in the network topology. This highlights the value of the MULTI approach, which preserves the structure and information inherent in individual layers, unlike UNFL. These findings also call into question the efficacy of the commonly used UNFL strategy, which is prevalent in the literature [19, 15, 34, 2, 31]. Results obtained with Infomap (reported in Appendix A) yield similar outcomes for IORussia. Conversely, for UK, UNFL shows only one lost community, whereas MULTI identifies several. This demonstrates that the choice of the community detection algorithm can significantly influence the results. In this case, the MULTI approach based on Generalized Infomap does not perform well, as it detects very large communities, collapsing nearly all nodes into a few oversized clusters.
@@
Union flattening. Figure 10a and 10c provide insights on the results of UNFL with respect to INDE. Hence, in UK dataset, UNFL exhibits similar patterns with blocks of common and gained communities and relatively few lost ones with respect to HST, MEN, and RTW. However, it shows a higher number of lost and gained communities compared with URL, suggesting that its information is poorly captured in the flattened network. Finally, UNFL has two communities in common with RPL and no lost communities but a significant number of gained ones, indicating that the detected communities differ substantially from those in the monomodal approach. Similar results are observed in IORussia: the Louvain approach shows a more stable situation, with only a few lost communities and many gained ones, whereas the Infomap-based approach loses a substantial number of communities compared to INDE. An exception emerges in UK, where UNFL loses only one community (with respect to the MEN layer) while detecting many common and gained communities relative to INDE.
@@
Multiplex community detection. Figure 10b and 10d demonstrate that MULTI has only two and one lost communities with respect to all monomodal approaches. For UK, the transition from the RPL layer to the multimodal approach reveals a higher number of gained communities, while other co-actions primarily show common communities and fewer gained ones. This suggests that communities identified using the multimodal MULTI approach retain much of the information from the monomodal approach while incorporating additional insights from the interplay between layers. Finally, in IORussia, the number of gained communties by MULTI approach is very high for all co-actions.
@@
Discussion. This analysis highlights the effectiveness of MULTI in preserving existing information while uncovering new communities. Meanwhile, the UNFL is more prone to losing communities, further underscoring its limitations. The results suggest that MULTI is more effective in capturing the underlying structure and information of the network, providing a more comprehensive view of coordinated behavior. The node-level analysis reveals similar trends, with MULTI outperforming the UNFL in retaining and identifying nodes. An exception arises in the UK dataset, where UNFL loses only one community and retrieves many common and gained ones, performing better than MULTI in this specific case. However, this remains an isolated outcome. Overall, the results consistently show that MULTI provides a more reliable and comprehensive view of coordinated behavior, while UNFL generally fails to capture important structural information. This comparison reinforces the importance of multimodal analysis for capturing the complexity of multiplex interactions, and underscores that the choice of community detection algorithm is equally crucial.
Figure 12: Lost, common, gained nodes characterization for UNFL and MULTI (Louvain) in UK dataset. We use the notation ns and * to indicate statistical significance levels. Specifically, ns denotes that there is no significant difference between the groups, with a p-value in the range $p\geq 0.05$. For statistically significant results, we used the following annotations based on p-value ranges: *: $p\leq 0.05$; **: $p\leq 0.01$; ***: $p\leq 0.001$.
@@ 5.2.2 Multimodal communities characterization
To assess the effectiveness of multimodal approaches in identifying central nodes, we analyze the lost, common, and gained nodes defined in Section 3.3.3 and shown in Figure 11. This evaluation aims to determine whether multimodal methods retain or lose key influential nodes. Losing peripheral nodes is not critical, but missing highly connected or bridging nodes would indicate a limitation. Conversely, successfully preserving or discovering new influential nodes would highlight the advantage of multimodality in capturing a more complete network structure. To conduct this analysis, we compute key centrality and structural metrics for each node including degree centrality, eigenvector centrality, local clustering coefficient, and PageRank. The comparison is performed for both UNFL and MULTI. Figure 12 illustrates, for UK with Louvain, the distribution of these metrics across lost, common, and gained nodes for each co-action. The results for IORussia and for Infomap approaches are reported in Appendix A for completeness.
We applied the Brunner-Munzel test to compare distributions across the lost, common, and gained node subsets for each metric. Unlike the Mann-Whitney U test, the Brunner-Munzel test does not assume equal variances or distribution shapes, making it more robust to ties, outliers, and variance differences, and thus more suitable for our analysis. The comparison highlights the performance of both the UNFL and MULTI, offering insights into their ability to retain or identify influential nodes.
@@
Union flattening. Common nodes are characterized by a higher degree centrality and eigenvector centrality in all combinations of datasets and approaches, with the exception of UK with Louvain, where lost and gained nodes exhibit higher values for HST, MEN, RTW, and URL. Instead, local clustering coefficient distribution is more homogeneous among the different co-actions, with some cases where common communities show higher values.
@@
Therefore, in case of HST, MEN, RTW, and URL, in UK dataset with Louvain, UNFL seems to lose the most important nodes, which are the most central and influential in the network, but at the same time it is able to detect new nodes that are not detected by the monomodal approach and that are also central and influential. In all the other cases, UNFL is able to retain the most important nodes in the network, which are the common ones, but at the same time it loses and gains nodes that are less central and influential.
@@
Multiplex community detection. In all combinations of datasets and approaches, observing MULTI, the common nodes have higher values of degree centrality and eigenvector centrality, with respect to the lost and gained nodes. There are some exceptions, such as in UK with Louvain, in which URL loses nodes that have higher values; or RPL in IORussia with Louvain, where gained nodes have higher centrality values. Local clustering coefficient shows more similar average for the lost, common, and gained nodes, but always with common nodes with higher values. Finally, page rank is on average similar for all the co-actions and for the lost, common, and gained nodes, but with a higher variance for the common nodes.
@@
Discussion. From the perspective of node centrality, both MULTI and UNFL approaches yield results that are generally consistent with those of the individual layers. Previous analyses showed that MULTI retains most of the key nodes and communities from each modality, and tends to detect additional ones, that are typically less central. The unique exception is in UK with Louvain, where UNFL exhibits the opposite behavior: it both gains and loses a notable number of nodes and communities, but crucially, it often discards highly central ones while still identifying new, influential actors. This suggests a significant misalignment between the results of UNFL and those of the individual layers.
5.3 Validation of coordinated behavior
Table 5: Validation of the detected communities in the IORussia coordination network using Louvain and Infomap community detection algorithms. For each layer and approach, we report the size of the largest three communities with size greater or equal to 70, along with the percentage of control and coordinated accounts in each community.
Louvain Infomap
approach size %ctrl %coord size %ctrl %coord
monomodal RTW 292 0.000 1.000 772 0.006 0.994
234 0.000 1.000 292 1.000 0.000
200 0.000 1.000 205 1.000 0.000
RPL 70 0.000 1.000 86 0.000 1.000
URL 208 0.000 1.000 337 1.000 0.000
182 1.000 0.000 214 0.000 1.000
82 1.000 0.000 115 0.000 1.000
MEN 324 0.012 0.988 $1,437$ 0.207 0.793
313 0.000 1.000 514 1.000 0.000
305 0.961 0.039 162 1.000 0.000
HST 318 0.000 1.000 $1,166$ 0.000 1.000
316 0.000 1.000 $559$ 0.998 0.002
217 0.995 0.005 $154$ 1.000 0.000
multimodal UNFL (nw) 476 1.000 0.000 $2,068$ 0.228 0.772
448 0.007 0.993 $1,761$ 0.999 0.001
442 1.000 0.000 $89$ 1.000 0.000
UNFL (ec) 485 1.000 0.000 $1,763$ 0.999 0.001
440 1.000 0.000 $1,588$ 0.002 0.998
421 0.971 0.029 $478$ 0.975 0.025
UNFL (sum) 475 1.000 0.000 $1,761$ 0.999 0.001
382 0.000 1.000 $1,588$ 0.002 0.998
370 0.968 0.032 $478$ 0.975 0.025
MULTI 1020 0.001 0.999 $5,012$ 0.194 0.806
922 1.000 0.000 $3,745$ 1.000 0.000
905 0.000 1.000 $145$ 1.000 0.000
The previous experiments focused primarily on assessing information that was lost, preserved, or gained across modalities, as well as on identifying signals of potential coordination. This section extends this analysis by directly validating the ability of the proposed methods to detect actual coordinated behavior, thereby strengthening the empirical support for our approach.
Validation in coordinated behavior detection is a crucial, yet challenging task due to the lack of ground truth data [1]. However, the IORussia dataset provides a unique opportunity for validation, as it includes labels distinguishing between coordinated and control accounts. Specifically, the IORussia dataset is composed by the coordinated group of accounts involved in the information operation and the control group of accounts. This labelling enables the validation of the coordinated communities detected by the different approaches. In Table 5, we report the validation of the detected communities in the IORussia coordination network. For INDE and multimodal approaches implemented with Louvain and Infomap, we report the size of the largest three communities, along with the percentage of control and coordinated accounts in each community. A good separation between these two groups indicates that the approach is effective in identifying coordinated behavior. Indeed, in all the approaches the communities are mainly composed by either the coordinated accounts or the control accounts, with a very small percentage from the other group. This confirms the effectiveness of all the approaches in identifying coordinated communities.
Figure 13: Number of coordinated communities lost, common, and gained detected by UNFL and MULTI in IORussia dataset.
Since the distinction between coordinated and non-coordinated communities is well defined, we cross-reference this information with the labels lost, common, and gained. Our goal is to assess how many coordinated communities are lost or gained by the UNFL and MULTI approaches. Figure 13 reports the number of coordinated communities detected by UNFL and MULTI (using both Louvain and Infomap). Specifically, Figure 13a shows the results for Louvain, where MULTI detects 40 coordinated communities – substantially more than UNFL. UNFL and MULTI detect, respectively, 12 and 15 coordinated common communities, showing similar detection power. Figures 13b shows the results for Infomap. Here, the number of coordinated communities are very similar across lost, common, and gained categories, with the only notable exception being that all three common communities detected by UNFL are coordinated. Overall, Louvain clearly performs better than Infomap, while MULTI is generally more effective at identifying a larger number of coordinated communities. We can conclude that the choice of community detection algorithm has a strong impact on the final outcomes.
Figure 14: Level of coordination of lost, common, and gained communities detected by Louvain approach.
Until now, we validated the results only on IORussia dataset since the labeling is available only for this dataset. However, we can also evaluate the results of the coordination detection through an internal measure, which is always available. Since the edge weights represent the level of coordination between two nodes, we can compute a coordination score for each community. Figure 14 and Figure A.13 in Appendix A show, respectively for Louvain and Infomap approach, the average weight, i.e. coordination level for each community grouped by the lost, common, gained labeling. We also report the mean of the averages for each label, indicating it with the horizontal line, colored according to the label.
In UK, lost and gained communities detected by MULTI exhibit higher coordination levels compared to those detected by UNFL, while common communities display similar coordination scores across the two approaches. This implies that, although MULTI loses only two communities, they are highly coordinated ones. However, Figure A.13A shows that UNFL loses several communities with high coordination scores as well as others with low scores, resulting in an overall mean that lies in the middle. In IORussia, the situation is more balanced across approaches. UNFL yields higher coordination levels for lost, common, and gained communities – both with Louvain and Infomap – compared to MULTI. Thus, in the IORussia case, the advantage of MULTI is less evident: the coordination levels of lost communities are not particularly low, and the coordination levels of gained communities are not especially high for either approach.
@@ 6 Discussion and conclusions
We explored the role of multimodality in the detection and analysis of coordinated behavior, addressing its challenges and opportunities. To tackle this complexity, we introduced a novel methodology that integrates multiple types of user actions into a unified analytical framework for detecting multimodal coordinated behavior. By comparing monomodal approaches with flattened representations and the multiplex community detection, our analysis highlighted the strengths and weaknesses of each approach, offering valuable insights into the detection and characterization of coordination in social media.
@@ 6.1 RQ1: Monomodal contributions
The second research question examines the contribution of multimodal approaches compared to monomodal ones in analyzing coordinated behavior. UNFL implemented with different flattening strategies produces similar outcomes, suggesting that the specific method of flattening has limited influence on the results (see heatmaps’ results in Appendix B,Appendix C, Appendix D, and Appendix E for empirical evidence). However, Section 5.2.1 demonstrates that the act of network flattening itself significantly impacts the detection of coordinated communities. In particular, the UNFL approach shares only a subset of communities with RTW, MEN, and HST, while failing to capture nearly all of those identified by URL. This indicates a substantial loss of information contributed by each monomodal layer. In contrast, the MULTI approach shows greater robustness: it retains almost all communities detected by monomodal methods and also uncovers additional coordinated structures. A notable exception emerges in the UK dataset with Infomap-based approaches, where UNFL loses only one community and identifies many common and gained ones, performing better than MULTI in this specific instance. However, this outcome remains isolated and does not generalize across datasets, but it highlights the importance of selection of the right community detection algorithm. Indeed, in literature Infomap is never used for coordinated behavior detection, while Louvain is the most exploited one [1]. These findings raise concerns about the use of network flattening in the literature and advocate for the adoption of more expressive, though complex, multimodal methods. Section 5.2.2 further supports the advantage of MULTI, highlighting its ability to preserve influential nodes while also identifying new ones. By comparison, also UNFL tends to detect the central nodes identified by INDE, but differently from MULTI it is not able, in general, to detect new communities. Overall, MULTI proves to be a more reliable approach for capturing both the breadth and depth of coordinated behavior.
@@ 6.3 Limitations and future works
This work presents the first comparative analysis of methods to operationalize the detection of multimodal coordinated behavior. Indeed, we defined multiple operationalizations of multimodality, assessing the contributions of individual monomodal approaches, and evaluating the benefits of a multimodal framework over monomodal ones. We conducted experiments on two datasets, UK and IORussia, comparing methods built on two community detection approaches, Infomap and Louvain. The primary limitation of this study is that the analysis was conducted only on Twitter/X platform, due to data availability. As a result, the generalizability to other social media platforms of our findings remains uncertain, and the conclusions drawn may be specific to the characteristics and dynamics of the observed cases. Future work will need to validate the proposed methodology across multiple platforms and datasets, ideally covering diverse social phenomena and contexts, to assess the robustness and consistency of our conclusions.
Our experiments focused on identifying information lost, preserved, or gained across modalities and on detecting signals of potential coordination. Section 5.3 complements this by validating the ability of the proposed methods to identify actual coordinated behavior. Nonetheless, these results are not universally generalizable: this remains a methodological study, and broader claims will require applying the approach in multiple empirical settings and events. Further validation in diverse real-world scenarios is therefore necessary to establish its wider applicability.
Such efforts will be crucial to determine whether the observed patterns hold more broadly and to establish the general applicability of our multimodal framework. Given the lack of prior literature on multimodal coordination, we addressed two foundational research questions, necessarily leaving out additional operationalizations of multimodality. For instance, exploring alternative community detection algorithms, such as multilayer clique percolation and ABACUS, could introduce new levels of multimodality within the intrinsic trade-off of coordination analysis (see Figure 2[38]. Furthermore, future work could refine the integration across layers by systematically investigating the $\omega$ and $w_{inter}$ parameters, which control the strength of inter-layer connections between the same user across modalities, respectively in Generalized Louvain and Generalized Infomap algorithms. Our current study focuses on five common co-actions, but the framework could be expanded to incorporate additional modalities, such as multimedia content sharing, which plays a central role on many social media platforms.
Moreover, although our approach identifies coordinated behavior through community structures, we acknowledge that real-world coordination may not always appear as clearly separated groups. By definition, coordination still involves multiple users acting synergistically, but these groups may overlap, with users participating in different coordinated activities at the same time. Overlapping community detection methods could capture such patterns [38], although they have not yet been explored in coordinated behavior research. Extending our framework to overlapping structures is an interesting future direction, particularly in the multimodal setting, where coordination signals are heterogeneous and the problem becomes even more challenging.